# Calibrated and Robust Foundation Models for Vision-Language and Medical Image Tasks Under Distribution Shift

### v1

@@ Abstract
Foundation models like CLIP and SAM have transformed computer vision and especially medical imaging via low-shot transfer learning, helping the cause of CADD with few case data. However, the deployment of these models hindered by two key challenges: distribution shift where training data differs from test data distribution (or pre-training from post-training distro) often due to inter-center image acquisition differences; and confidence misalignment that leads to overconfident incorrect predictions. These issues manifest differently in vision-language classification and medical segmentation tasks (how differently, and how does domain-specificity arise/?), yet existing solutions remain domain-specific. We propose a unified framework to address these two challenges together, named StaRFM. It introduces a Fisher information penalty (FIP), extended to $3$D medical data via patch-wise regularization, to reduce covariate shift in CLIP and SAM embeddings. Additionally, a confidence misalignment penalty (CMP), reformulated for voxel-level predictions, calibrates uncertainty in segmentation tasks. We theoretically derive PAC-Bayes bounds showing FIP controls generalization via the Fisher-Rao norm, while CMP minimizes calibration error through Brier score optimization. StaRFM outperforms all baselines by +3.5% accuracy and 28% lower ECE on 19 vision datasets (e.g., ImageNet, Office-Home), 84.7% DSC and 4.8mm HD95 in medical segmentation (e.g., BraTS, ATLAS), and 40% lower cross-domain performance gap compared to prior benchmarking methods. The framework is plug-and-play, requiring minimal architectural changes for integration with foundation models. Source code and models will be released at the anonymous link: StaRFM.
* Corresponding author
Keywords Foundation models  $\cdot$ Prompt learning  $\cdot$ Confidence Calibration  $\cdot$ Segment Anything Model
@@ 1 Introduction
Recent advances in foundation models such as CLIP Radford et al. (2021), Align Abdul Samadh et al. (2024) and Flamingo Alayrac et al. (2022), in vision-language and Segment Anything Model Kirillov et al. (2023), MedSAM Ma et al. (2024), BioMedCLIP Zhang et al. (2023), UNET Ronneberger et al. (2015), Dinov2 Oquab et al. (2023), and CXR-CLIP You et al. (2023) in medical image analysis have revolutionized machine learning by enabling zero-shot transfer learning to downstream tasks. However, their real-world deployment is hindered by two pervasive intertwined challenges: Distribution shift where test data diverge from training distribution ($P_{tr}(y,x)\neq P_{ts}(y,x)$), and confidence misalignment, where models produce overconfident yet incorrect predictions Guo et al. (2017); Pereyra et al. (2017). These issues are particularly acute in safety-critical domains, from autonomous systems to medical diagnostics, where unreliable uncertainty estimates can lead to catastrophic failures Yao et al. (2024).
In vision-language models (VLMs) like CLIP, Covariate shift where the source and target feature distributions differs ($P_{tr}(x)\neq P_{ts}(x)$) Shimodaira (2000); Sugiyama et al. (2007) while $P(y|x)$ remains unchanged. This shift is ubiquitous in real-world settings such as cross-domain visual classification, population-biased medical imaging, or protocol variations across clinical institutions Khan et al. (2025). Concurrently, such shifts exacerbate confidence misalignment, where softmax probabilities fail to reflect true likelihoods Murugesan et al. (2025).
Recent work Khan et al. (2025) proposed CalShift, a unified framework addressing both challenges via Fisher information regularization and a confidence misalignment penalty (CMP). While effective for 2D classification, its applicability to high-dimensional, spatially structured data (e.g., 3D medical images) remains unexplored.
@@
(b) Aligned predictions after integration of CMP and FIM $I(\theta)$.
Figure 1: Workflow of the proposed StaRFM framework: The sub-figure (a) illustrates the confidence misalignment problem caused by covariate shift. The (bottom right) part of the sub-figure (a) show misaligned predictions. The subfigure (b) middle section represents the two components as recipe in method: Fisher information penalty ($I(\theta)$) for covariate shift correction and confidence misalignment penalty (CMP) for calibration. Both FIM and CMP are integrated into the CLIP text encoder. The final loss, $\mathcal{L}_{\text{StaRFM}}=\mathcal{L}_{\text{CLIP}}+\lambda_{1}I(\theta)+\lambda_{2}\text{CMP}$, combines the original CLIP loss with FIM and CMP to produce robust and aligned predictions, as shown in (bottom right) as output.
@@
    Through extensive experiments on 19 vision datasets and 3 medical benchmarks (BraTS, ATLAS), we show that our method is consistent in improving performance over different metrics ?, achieving up to +3.5% accuracy and 28% lower Expected Calibration Error (ECE) compared to baselines.
@@
    Our method operates as a plug-and-play module for the text encoder used with foundation models such as CLIP and SAM, requiring minimal architectural changes and training effort.
@@
Figure 2: Aligned predictions after integration of CMP and FIM $I(\theta)$.
Figure 3: Workflow of the proposed StaRFM framework for medical segmentation tasks. Both FIM and CMP are integrated into the SAM prompt encoder. The final loss, $\mathcal{L}_{\text{StaRFM}}=\mathcal{L}_{\text{SAM}}+\lambda_{1}I_{3D}(\theta)+\lambda_{2}\text{CMP}_{3D}$, combines the original SAM loss with FIM and CMP to produce robust and aligned predictions.
@@ 2 Methods in comparison
Foundation models in vision and medicine. Foundation models like CLIP Radford et al. (2021), Flamingo Alayrac et al. (2022), SAM Kirillov et al. (2023), and MedSAM Ma et al. (2024) have transformed computer vision and clinical applications via zero-shot transfer. CLIP’s contrastive image-text pretraining enables strong downstream classification Zhou et al. (2022), while SAM’s prompt-based segmentation generalizes well on natural images. However, medical adaptations (e.g., MedSAM Ma et al. (2024), BioMedCLIP Zhang et al. (2023)) face challenges under distribution shifts. Such as, covariate shift arises when CLIP’s pretrained features ($P_{tr}(x)$) diverge from target distributions ($P_{ts}(x)$Khan et al. (2024), while in medical imaging SAM yields overconfident errors on MRI/CT scans due to domain gaps with natural images Huang et al. (2024).
Covariate shift in vision-language and medical tasks. Covariate shift appears differently in vision-language models (VLMs) and medical image segmentation. In CLIP-like architectures, it occurs when pretrained text-image feature distributions ($P_{src}(x)$) diverge from downstream task data ($P_{tgt}(x)$Khan et al. (2024). This divergence often results from dataset fragmentation in few-shot learning Zhou et al. (2022) and domain gaps between web-scale pretraining (e.g., LAION) and specific target domains Abdul Samadh et al. (2024). While prior work quantifies this shift using Fisher Information $I(\theta)$ Khan and Syed (2025), it overlooks the complexities of 3D data.
@@
We introduce a unified Confidence misalignment penalty (CMP) by extending Khan and Syed (2025)’s formulation to voxel-level predictions which is given in eq. 6.
@@ 3.1 Problem Formulation
Vision-Language Tasks. The input to vision-language tasks consists of an image $x\in\mathbb{R}^{H\times W\times 3}$ and a text prompt $t$. The output is a set of class probabilities $P(y|x,t)\in\mathbb{R}^{K}$, which is obtained through the softmax function. The objective is to minimize the expected loss, defined as
@@
Fisher Information Penalty (FIP) The Fisher Information Penalty (FIP) is a metric that measures the sensitivity of model parameters to distribution shifts. It is defined as the expected value of the negative Hessian of the log-likelihood function, as shown in the following equation:
@@
Task-Specific Implementations: In vision tasks, such as in the CLIP model, the FIP is computed globally over image-text embeddings. The resulting computation is given by:
@@
Confidence Misalignment Penalty (CMP) The Confidence Misalignment Penalty (CMP) is designed to penalize overconfident incorrect predictions. It is mathematically defined as:
@@
Task-specific adaptations: For vision tasks, such as image classification, the CMP is applied at the class level, where the penalty is computed over the predicted class probabilities for the image $x_{\text{img}}$ and target class $t$:
@@
Task-specific loss functions
@@
Proposition 3.1 (Fisher Regularization for Domain Adaptation).
@@
$\mathcal{R}_{\text{tgt}}(\theta)\leq\mathcal{R}_{\text{src}}(\theta)+\frac{1}{2}\sqrt{\mathrm{tr}(I(\theta)\Sigma_{\text{shift}})}+\mathcal{O}(n^{-1/2})$ (11)
@@
where $\Sigma_{\ell}$ denotes the covariance of loss gradients. Regularization via $\|I(\theta)\|_{*}$ minimizes this bound:
$\min_{\theta}\mathcal{R}_{\text{src}}(\theta)+\lambda\|I(\theta)\|_{*}\quad\Rightarrow\quad(I(\theta)\Sigma_{\text{shift}})\leq\epsilon$ (12)
@@
Loss in medical image segmentation.
Our segmentation loss combines region-based and voxel-wise measures with distribution-aware regularization. Dice Loss Sudre et al. (2017): Measures volumetric overlap between predicted mask $\mathbf{M}$ and ground truth $\mathbf{M}^{*}$:
@@ 4 Experiments
We evaluate our method across two distinct domains. Vision-language tasks using 19 standard vision benchmarking datasets (including vision and domain adaptation), and 3D medical image segmentation with multi-institutional brain MRI collections such as BRATS and ATLAS. Our experiments provide empirical support to three core claims:
@@ 4.1 Dataset and Tasks
Medical image segmentation tasks. We use three publicly available brain MRI datasets in our experiments to evaluate the effectiveness of our method in medical image segmentation tasks. BraTS 2023Menze et al. (2014) includes 1,250 multi-parametric MRI scans (T1, T1c, T2, FLAIR) collected from 60 centers, annotated with tumor subregions. ATLAS v2.0Liew et al. (2022) comprises 1,128 T1-weighted scans from 22 international sites, each annotated with chronic stroke lesions. IBSR18 consists of legacy 1.5T T1-weighted scans and is used to evaluate robustness to low-resolution data.
@@ 4.3 Evaluation Metrics
We use task-specific metrics for performance assessment of our method. For vision tasks, we report classification accuracy as robustness metric in vision-language tasks. For medical segmentation, we use Dice Similarity Coefficient (DSC) and 95% Hausdorff Distance (HD95).
Calibration. is measured using expected calibration rrror (ECE) Guo et al. (2017), computed as:
@@
where $S_{b}$ denotes the set of samples in bin $b$, with $B=10$ bins.
Domain robustness. is quantified using two metrics: Domain Generalization Gap (DGG), defined as $\text{acc}_{\text{src}}-\text{acc}_{\text{tgt}}$, and Cross-Site Variance, computed as the standard deviation of DSC across different institutions.
@@ 5 Results and Discussion
FIP yields the most significant gains in the 0-shot to 4-shot range (e.g., +5.1% for CoOp at 0-shot, +3.5% at 2-shot), where limited training data exacerbates covariate shift. This aligns with our theoretical analysis (Proposition 1 LABEL:), as FIM’s curvature-based regularization helps mitigate overfitting when samples are scarce.
As the number of shots increases, performance gains diminish—shrinking to $\leq 0$ in the 16-shot setting—which suggests that FIM’s primary role is to stabilize feature distributions under data scarcity. For instance, MaPLe+FIM shows marginal improvement (+0.8% at 2-shot) and slight degradation (-1.0% at 8-shot), indicating that well-regularized baselines (e.g., MaPLe’s multi-modal alignment) benefit less from additional regularization.
@@
While CMP independently improves calibration, combining it with FIM yields Pareto-optimal gains in both accuracy (Table 2) and ECE. For example, CoOp+FIM+CMP achieves 12.5$\downarrow$ ECE vs. 8.05$\downarrow$ with CMP alone.
Table 2: Our method few-shot accuracy performance results on ImageNet dataset with and without FIM penalty. The $\Delta$ row shows the percentage increase ($\uparrow$) or decrease ($\downarrow$) in accuracy.
@@
Table 3: Our method few-shot ECE performance results on ImageNet dataset with and without CMP penalty. The $\Delta$ row shows the percentage increase ($\uparrow$) decrease ($\downarrow$) in calibration
@@
StaRFM accuracy performance on vision datasets. The comprehensive evaluation in Table 4 demonstrates the effectiveness of our method StaRFM across diverse vision tasks, validating its dual approach of Fisher Information Matrix (FIM) regularization for accuracy improvement and Confidence Misalignment Penalty (CMP) for calibration enhancement.
FIP regularization provides accuracy improvements across nearly all datasets, with an average gain of 3.2%. The most significant improvements occur in action recognition (UCF101: +7.3%) and fine-grained classification (Flowers102: +8.0%), where domain shifts are particularly challenging. While most datasets benefit from FIP, we observe diminishing returns on already well-performing benchmarks (Caltech101: +0.2%) and slight degradation on StanfordCars (-8.2%). This aligns with our theoretical analysis showing FIM is most beneficial when significant covariate shift exists between training and test distributions. The improvements on ImageNet (+6.8%) and its derivatives (Food101: +1.7%) demonstrate FIM’s effectiveness in handling real-world distribution shifts, a finding consistent with our few-shot experiments in Table 2.
StaRFM calibration performance on vision datasets. CMP reduces Expected Calibration Error (ECE) by an average of 5.7%, with particularly strong improvements on Food101 (-9.85%) and OxfordPets (-7.19%). This validates CMP’s mechanism for redistributing probability mass from overconfident incorrect predictions.
The Flowers102 dataset shows a 6.76% increase in ECE, suggesting that CMP may over-penalize in cases where the label distribution is naturally imbalanced.
The simultaneous improvement in both accuracy (upper section) and calibration (lower section) demonstrates that FIM and CMP address orthogonal challenges in model deployment, supporting their joint use in the complete StaRFM framework.
@@
Table 4: The upper half shows StaRFM accuracy results on vision datasets with and without FIM penalty. The $\Delta$ row shows the percentage increase ($\uparrow$) or decrease ($\downarrow$) in accuracy. The lower half shows StaRFM ECE results on vision datasets with and without CMP penalty. The $\Delta$ row shows the percentage increase ($\uparrow$) decrease ($\downarrow$) in calibration
@@
Table 5: The upper part shows our method accuracy results on covariate shift vision datasets with and without FIM penalty. The $\Delta$ row shows the percentage increase/decrease in accuracy. $\uparrow$ shows improvement in accuracy while the $\downarrow$ shows decrease in performance. The lower part shows our method ECE results on covariate shift vision datasets with and without CMP penalty. The $\Delta$ row shows the percentage increase/decrease in calibration. $\downarrow$ shows improvement in calibration while $\uparrow$ shows calibration performance decrease.
@@
StaRFM performance across modalities. Table LABEL:tab:bmTbl highlights StaRFM’s strong performance in clinical segmentation tasks, extending the improvements shown in our vision-language results (Table 4).

### v2

@@ Abstract
Foundation models like CLIP and SAM have advanced computer vision and medical imaging via low-shot transfer learning, aiding CADD with limited data. However, their deployment faces two key challenges. distribution shift where pre-training and post-training data distributions differ (e.g., due to inter-center image acquisition) and confidence misalignment, which leads to overconfident errors. These issues surface differently, vision-language models (e.g., CLIP) suffer from 2D embedding shift (image-text misalignment), while medical models (e.g., SAM) encounter 3D domain shifts (e.g., scanner variation) and voxel-wise calibration need. Existing solutions are domain-specific. We propose StaRFM, a fusion of Fisher information penalty (FIP) and confidence misalignment penalty (CMP) tackling both challenges. It applies FIP, extended to 3D via patch-wise regularization, to reduce embedding shift, and CMP, reformulated for voxel-level predictions, to calibrate segmentation uncertainty. We derive PAC-Bayes bounds. FIP controls generalization via the Fisher-Rao norm, and CMP reduces calibration error via Brier score minimization. StaRFM surpasses baselines by +3.5% accuracy and 28% lower ECE on 19 vision datasets (e.g., ImageNet, Office-Home), achieves +4.2% DSC over SAM-FT and 4.8mm HD95 on medical benchmarks (e.g., BraTS, ATLAS), and reduces cross-domain gaps by up to 20%. The framework is plug-and-play, requiring minimal architectural changes. Code and models are available at: StaRFM.
keywords:
[…]
@@ 1 Introduction
Recent advances in foundation models such as CLIP Radford et al. (2021), Align Abdul Samadh et al. (2024) and Flamingo Alayrac et al. (2022), in vision-language and Segment Anything Model Kirillov et al. (2023), MedSAM Ma et al. (2024), BioMedCLIP Zhang et al. (2023), UNET Ronneberger et al. (2015), Dinov2 Oquab et al. (2023), and CXR-CLIP You et al. (2023) in medical image analysis have revolutionized machine learning by enabling zero-shot transfer learning to downstream tasks. However, their real-world deployment is hindered by two pervasive intertwined challenges. Distribution shift where test data diverge from training distribution $P_{tr}(y,x)\neq P_{ts}(y,x)$, and confidence misalignment, where models produce overconfident yet incorrect predictions Guo et al. (2017); Pereyra et al. (2017). These issues are particularly acute in safety-critical domains, from autonomous systems to medical diagnostics, where unreliable uncertainty estimates can lead to catastrophic failures Yao et al. (2024); Mirza et al. (2021).
In vision-language models (VLMs) like CLIP, Covariate shift where the source and target feature distributions differ ($P_{tr}(x)\neq P_{ts}(x)$) Shimodaira (2000); Sugiyama et al. (2007) while $P(y|x)$ remains unchanged. This shift is ubiquitous in real-world settings such as cross-domain visual classification, population-biased medical imaging, or protocol variations across clinical institutions Khan et al. (2025). Concurrently, such shifts exacerbate confidence misalignment, where softmax probabilities fail to reflect true likelihoods Murugesan et al. (2025).
Recent work Khan et al. (2025) proposed CalShift, a unified framework addressing both challenges via Fisher information regularization and a confidence misalignment penalty (CMP). While effective for 2D classification, its applicability to high-dimensional, spatially structured data (e.g., 3D medical images) remains unexplored. In this work, we extend CalShift Khan et al. (2025) published in CVPR2025, from 2D classification to vision-language and 3D medical imaging tasks.
@@
(b) Aligned predictions after integration of CMP and FIP $I(\theta)$.
Figure 1: Workflow of the proposed StaRFM framework: The sub-figure (a) illustrates the confidence misalignment problem caused by covariate shift. The (bottom right) part of the sub-figure (a) show misaligned predictions. The subfigure (b) middle section represents the two components as recipe in method: Fisher information penalty ($I(\theta)$) for covariate shift correction and confidence misalignment penalty (CMP) for calibration. Both FIP and CMP are integrated into the CLIP text encoder. The final loss, $\mathcal{L}_{\text{StaRFM}}=\mathcal{L}_{\text{CLIP}}+\lambda_{1}I(\theta)+\lambda_{2}\text{CMP}$, combines the original CLIP loss with FIP and CMP to produce robust and aligned predictions, as shown in (bottom right) as output.
@@
    Through extensive experiments on 19 vision datasets , 8 domain shift and 3 medical benchmarks (BraTS, ATLAS), we demonstrate that our method consistently improves performance metrics such as accuracy, DSC, and HD95, achieving up to a +3.5% increase in accuracy and a 28% reduction in Expected Calibration Error (ECE) compared to the baselines.
@@
    StaRFM operates as a plug-and-play module for the text encoder used with foundation models such as CLIP and SAM, requiring minimal architectural changes and training effort.
@@
Figure 2: Workflow of the proposed StaRFM framework for medical segmentation tasks. Both FIP and CMP are integrated into the SAM prompt encoder. The final loss, $\mathcal{L}_{\text{StaRFM}}=\mathcal{L}_{\text{SAM}}+\lambda_{1}I_{3D}(\theta)+\lambda_{2}\text{CMP}_{3D}$, combines the original SAM loss with FIP and CMP to produce robust and aligned predictions.
@@ 2 Methods in comparison
Foundation models in vision and medicine. Foundation models like CLIP Radford et al. (2021), Flamingo Alayrac et al. (2022), SAM Kirillov et al. (2023), and MedSAM Ma et al. (2024) have transformed computer vision and clinical applications via zero-shot transfer. CLIP’s contrastive image-text pretraining enables strong downstream classification Zhou et al. (2022), while SAM’s prompt-based segmentation generalizes well on natural images. However, medical adaptations (e.g., MedSAM Ma et al. (2024), BioMedCLIP Zhang et al. (2023)) face challenges under distribution shifts. Such as, covariate shift arises when CLIP’s pretrained features $P_{tr}(x)$ diverge from target distributions $P_{ts}(x)$ Khan et al. (2024), while in medical imaging SAM exhibits overconfident errors on MRI and CT scans due to domain gaps with natural images Huang et al. (2024).
Covariate shift in vision-language and medical tasks. Covariate shift appears differently in vision-language models (VLMs) and medical image segmentation. In CLIP-like architectures, it occurs when pretrained text-image feature distributions $P_{src}(x)$ diverge from downstream task data $P_{tgt}(x)$ Khan et al. (2024). This divergence often results from dataset fragmentation in few-shot learning Zhou et al. (2022) and domain gaps between web-scale pretraining (e.g., LAION) and specific target domains Abdul Samadh et al. (2024). While prior work quantifies this shift using Fisher information $I(\theta)$ Khan and Syed (2025), it overlooks the complexities of 3D data.
@@
We introduce StaRFM which unify CMP by extending Khan and Syed (2025) formulation to voxel-level predictions, and Fisher information Khan et al. (2024) to 3D data via patch-wise gradients, as given in Eq. 6.
@@ 3.1 Problem Formulation
Vision-Language Tasks. The input to vision-language tasks consists of an image $x\in\mathbb{R}^{H\times W\times 3}$ and a text prompt $t$. The output is a set of class probability $P(y|x,t)\in\mathbb{R}^{K}$, which is obtained through the softmax function. The objective is to minimize the expected loss, defined as
@@
Fisher Information Penalty (FIP). The FIP is a metric that measures the sensitivity of model parameters to distribution shifts. It is defined as the expected value of the negative Hessian of the log-likelihood function, as shown in the following equation:
@@
Task-Specific Implementations. In vision tasks, such as in the CLIP model, the FIP is computed globally over image-text embeddings. The resulting computation is given by:
@@
Confidence Misalignment Penalty. The Confidence Misalignment Penalty (CMP) is designed to penalize overconfident incorrect predictions. It is mathematically defined as:
@@
Task-specific adaptations. For vision tasks, such as image classification, the CMP is applied at the class level, where the penalty is computed over the predicted class probabilities for the image $x_{\text{img}}$ and target class $t$:
@@
3.2 Task-specific loss functions
@@
Proposition 3.1.
@@
$\mathcal{R}_{\text{tgt}}(\theta)\leq\mathcal{R}_{\text{src}}(\theta)+\frac{1}{2}\sqrt{(I(\theta)\Sigma_{\text{shift}})}+\mathcal{O}(n^{-1/2})$ (11)
@@
where $\Sigma_{\ell}$ denotes the covariance of loss gradients. Regularization via $\|I(\theta)\|$ minimizes this bound:
$\min_{\theta}\mathcal{R}_{\text{src}}(\theta)+\lambda\|I(\theta)\|\quad\Rightarrow\quad(I(\theta)\Sigma_{\text{shift}})\leq\epsilon$ (12)
@@
Theorem 3.1.
Let $\mathcal{D}_{\text{src}},\mathcal{D}_{\text{tgt}}$ be source and target distributions under covariate shift such that $P_{\text{src}}(x)\neq P_{\text{tgt}}(x)$ but $P_{\text{src}}(y|x)=P_{\text{tgt}}(y|x)$. For a segmentation model $f_{\theta}:x\mapsto\hat{y}$ with voxel-level probabilistic predictions $P(y_{v}|x_{v})$, define the confidence misalignment penalty (CMP) as:
$\text{CMP}_{3D}=\frac{1}{|\mathcal{V}|}\sum_{v\in\mathcal{V}}\frac{P(y_{v}^{\prime}|x_{v})}{1-P(y_{v}^{\prime}|x_{v})},$
where $y_{v}^{\prime}=\arg\max_{k}P(y_{v}=k|x_{v})$ denotes the predicted label at voxel $v$.
Then, for any $\delta\in(0,1)$ and with probability at least $1-\delta$, the expected calibration error (ECE) is bounded as:
$\text{ECE}(f_{\theta})\leq\sqrt{\frac{\text{CMP}_{3D}}{n}}+\underbrace{\mathbb{E}_{x,y}\left[(y_{v}-P(y_{v}|x_{v}))^{2}\right]}_{\text{Brier Score}}.$
Proof.
Following the decomposition of calibration error from Kumar et al. (2019), the ECE over voxels can be defined as:
$\text{ECE}(f_{\theta})=\frac{1}{|\mathcal{V}|}\sum_{v\in\mathcal{V}}\left|P(y_{v}|x_{v})-\mathbb{E}[y_{v}|x_{v}]\right|.$
From Guo et al. (2017), we know that:
$\left|P(y_{v}|x_{v})-\mathbb{E}[y_{v}|x_{v}]\right|\leq\sqrt{\mathbb{E}[(y_{v}-P(y_{v}|x_{v}))^{2}]}=\sqrt{B_{v}},$
where $B_{v}$ is the voxel-wise Brier score. Under confidence misalignment, overconfident predictions dominate the calibration gap, particularly when $P(y_{v}^{\prime}|x_{v})\gg P(y_{v}|x_{v})$. Thus, we bound:
$B_{v}\leq\text{CMP}_{3D}^{(v)}=\frac{P(y_{v}^{\prime}|x_{v})}{1-P(y_{v}^{\prime}|x_{v})}.$
Aggregating over $n$ samples and applying Hoeffding’s inequality:
$\text{ECE}(f_{\theta})\leq\sqrt{\frac{\text{CMP}_{3D}}{n}}+\mathbb{E}_{x,y}\left[(y_{v}-P(y_{v}|x_{v}))^{2}\right].$
Loss in medical image segmentation. Our segmentation loss combines region-based and voxel-wise measures with distribution-aware regularization. Dice Loss Sudre et al. (2017) Measures volumetric overlap between predicted mask $\mathbf{M}$ and ground truth $\mathbf{M}^{*}$:
@@ 4 Experiments
We evaluate our method across two distinct domains. Vision-language tasks using 19 standard vision benchmarking datasets (including vision and domain adaptation), and 3D medical image segmentation with multi-institutional brain MRI collections such as BRATS and ATLAS. Our experiments provides empirical support to three core claims:
@@ 4.1 Dataset and Tasks
Medical image segmentation tasks. We use three publicly available brain MRI datasets in our experiments to evaluate the effectiveness of our method in medical image segmentation tasks. BraTS 2023Menze et al. (2014) includes 1,250 multi-parametric MRI scans (T1, T1c, T2, FLAIR) collected from 60 centers, annotated with tumor subregions. ATLAS v2.0 Liew et al. (2022) comprise 1,128 T1-weighted scans from 22 international sites, each annotated with chronic stroke lesions.
@@ 4.3 Evaluation Metrics
We use domain-specific metrics to evaluate StaRFM performance. For vision-language tasks, we report classification accuracy following standard protocol in CLIP-based adaptation studies Radford et al. (2021); Zhou et al. (2022), measuring the proportion of correctly predicted labels. For medical segmentation, we adopt the clinically established Dice Similarity Coefficient (DSC) Sudre et al. (2017) for volumetric overlap and 95% Hausdorff Distance (HD95) Karimi and Salcudean (2019) for boundary alignment.
Calibration. We quantify prediction reliability using Expected Calibration Error (ECE) Guo et al. (2017); Naeini et al. (2015), which measures the discrepancy between model confidence and empirical accuracy across probability bins:
@@
where $S_{b}$ denotes samples in confidence bin $b$ ($B=10$), $N$ is total samples, acc(·) is bin accuracy, and conf(·) is mean predicted confidence.
Domain robustness. We assess cross-domain stability through two metrics: Domain Generalization Gap (DGG) Gulrajani and Lopez-Paz (2020), calculated as $\text{acc}_{\text{src}}-\text{acc}_{\text{tgt}}$, and Cross-Site Variance Menze et al. (2014), computed as the standard deviation of DSC across institutional test sets.
@@ 5 Results and Discussion
FIP exhibits the most significant gains in the 0-shot to 4-shot range (e.g., +5.1% for CoOp at 0-shot, +3.5% at 2-shot), where limited training data exacerbates covariate shift. This aligns with our theoretical analysis (Proposition 1 3.1), as FIP curvature-based regularization helps mitigate overfitting when samples are scarce.
As the number of shots increases, performance gains diminish shrinking to $\leq 0$ in the 16-shot setting which suggests that FIP primary role is to stabilize feature distributions under data scarcity. For instance, MaPLe+FIP shows marginal improvement (+0.8% at 2-shot) and slight degradation (-1.0% at 8-shot), indicating that well-regularized baselines (e.g., MaPLe’s multi-modal alignment) benefit less from additional regularization.
@@
While CMP independently improves calibration, combining it with FIP manifest pareto-optimal gains in both accuracy (Table 2) and ECE. For example, CoOp+FIP+CMP achieves 12.5$\downarrow$ ECE vs. 8.05$\downarrow$ with CMP alone.
Table 2: StaRFM few-shot accuracy performance results on ImageNet dataset with and without FIP penalty. The delta ($\Delta$) row shows the percentage increase ($\uparrow$) or decrease ($\downarrow$) in accuracy.
@@
Table 3: StaRFM few-shot ECE performance results on ImageNet dataset with and without CMP penalty. The delta ($\Delta$) row shows the percentage increase ($\uparrow$) or decreases ($\downarrow$) in calibration
@@
StaRFM accuracy performance on vision datasets. The comprehensive evaluation in Table 4 demonstrates the effectiveness of our method StaRFM across diverse vision tasks, validating its dual approach of FIP regularization for accuracy improvement and CMP for calibration enhancement.
FIP regularization provides accuracy improvements across nearly all datasets, with an average gain of 3.2%. The most significant improvements occur in action recognition (UCF101: +7.3%) and fine-grained classification (Flowers102: +8.0%), where domain shifts are particularly challenging. While most datasets benefit from FIP, we observe diminishing returns on already well-performing benchmarks (Caltech101: +0.2%) and slight degradation on StanfordCars (-8.2%). This aligns with our theoretical analysis showing FIP is most beneficial when significant covariate shift exists between training and test distributions. The improvements on ImageNet (+6.8%) and its derivatives (Food101: +1.7%) demonstrate FIP’s effectiveness in handling real-world distribution shifts, a finding consistent with our few-shot experiments in Table 2.
StaRFM calibration performance on vision datasets. CMP reduces ECE by an average of 5.7%, with particularly strong improvements on Food101 (-9.85%) and OxfordPets (-7.19%). This validates CMP’s mechanism for redistributing probability mass from overconfident incorrect predictions. The Flowers102 dataset shows a 6.76% increase in ECE, suggesting that CMP may over-penalize in cases where the label distribution is naturally imbalanced. The simultaneous improvement in both accuracy (upper section) and calibration (lower section) demonstrates that FIP and CMP address orthogonal challenges in model deployment, supporting their joint use in the complete StaRFM framework.
@@
Table 4: The upper half shows StaRFM accuracy results on vision datasets with and without FIP penalty. The delta ($\Delta$) row shows the percentage increase ($\uparrow$) or decrease ($\downarrow$) in accuracy. The lower half shows StaRFM ECE results on vision datasets with and without CMP penalty. The delta ($\Delta$) row shows the percentage increase ($\uparrow$) decrease ($\downarrow$) in calibration
@@
Table 5: The upper part shows StaRFM accuracy results on covariate shift vision datasets with and without FIP penalty. The delta ($\Delta$) row shows the percentage increase/decrease in accuracy. $\uparrow$ shows improvement in accuracy while the $\downarrow$ shows decrease in performance. The lower part shows StaRFM ECE results on covariate shift vision datasets with and without CMP penalty. The $\Delta$ row shows the percentage increase/decrease in calibration. $\downarrow$ shows improvement in calibration while $\uparrow$ shows calibration performance decrease.
@@
StaRFM performance across modalities. Table 6 highlights StaRFM’s strong performance in clinical segmentation tasks, extending the improvements shown in our vision-language results (Table 4).